OpenStudy (saraa23):

I am struggling with double angle formulas and inverse trig. How would you rewrite the following as an algebraic expression of x. Cos(2tan^-1(x))

OpenStudy (mathmale):

Please present "inverse trig functions" by drawing them in the Draw utility or through Equation Editor. \[\cos (\tan ^{-1}x)\]

OpenStudy (mathmale):

This reads: "Find the cosine of the angle whose tangent is x." I could and will expand that: "Find the cosine of the angle whose tangent is x/1."

OpenStudy (mathmale):

Can you draw an angle that has a tangent of x/1? Here the opposite side is represented by x and the adjacent side is represented by 1. Use the Draw utility, below.

OpenStudy (mathmale):



OpenStudy (saraa23):

my picture is not wanting to send @mathmale. Sorry if I'm no typing the equations as they're supposed to be I am new to this website.

OpenStudy (jdoe0001):

what's the tangent identity equals to anyway? I mean hmm \(\bf sin(\theta)=\cfrac{opposite}{hypotenuse} \qquad \qquad % cosine cos(\theta)=\cfrac{adjacent}{hypotenuse} \\ \quad \\ % tangent tan(\theta)=?\)

OpenStudy (saraa23):

it is equal to OPPOSITE/ADJACENT

OpenStudy (jdoe0001):

ok... so hmmm \(\bf tan^{-1}(x)=\theta \qquad \textit{really means}\qquad tan(\theta)=x\implies tan(\theta)=\cfrac{x}{1}\) maybe?

OpenStudy (jdoe0001):

what... hmm do you think?

OpenStudy (saraa23):

Okay what you did right there is you basically interchanged the y and the x to make f'(x) into the original f(x). So does that break the equation down to \[Cos(2(x/1))\]

OpenStudy (saraa23):

@jdoe0001

OpenStudy (phi):

let A be an angle then you want cos(2A) A= tan^-1 (x) so tan A = x/1 draw a triangle, label the opposite and advacent sides (x and 1) find the hypotenuse then you can find cos A and sin A and use those to find cos(2A)

OpenStudy (jdoe0001):

well... x/1 is the tangent value, not the angle itself, the inverse tangent function, is spitting out the angle itself so..... now that you know who the opposite and trhe adjacent sides are, graph them as mathmale suggested ::)

OpenStudy (jdoe0001):

notice the angle, say \(\bf tan^{-1}(x)={\color{brown}{ \theta }}\qquad \textit{really means}\qquad tan(\theta)=x\implies tan(\theta)=\cfrac{x}{1} \\ \quad \\ thus \\ \quad \\ cos(2tan^{-1}(x))\implies cos(2{\color{brown}{ \theta}})\implies 2cos^2({\color{brown}{ \theta}})-1\implies ?\) so, if you just graph it, and from there, you'll just need the cosine of it

OpenStudy (saraa23):

after finding the hypo. I come up with \[Cos(\cos ^{2}(1\div \sqrt{x ^{2}+1}) - \sin ^{2}(x \div \sqrt{x ^{2}+1}))\] is that correct ? @jdoe0001

OpenStudy (phi):

you might want to think like this cos(2A) = cos^2 A - sin^2 A you know cos A and sin A so you can find cos(2A)

OpenStudy (phi):

It can get confusing, so I try to keep things simple I know tan^-1 (x) is some angle, so call it A from which we get tan(A)= x/1 from that we get sin(A) x/sqr(x^2+1) and cos A 1/sqr(x^2+1) we want cos(2A) (replace tan^-1 x with the name of the angle A) cos(2A) by the double angle formula is cos^2 A - sin^2 A plug in cosA and sinA and we get our answer

OpenStudy (phi):

not clear ?

OpenStudy (saraa23):

when putting into the double angle formula A is x/1 ?

OpenStudy (phi):

no. The logic is not quite "plugging in" things. We are being tricky, and have to think a bit differently. Start with this: if you want cos(2A) and I tell you cosA and sin A , can you find cos2A ?

OpenStudy (saraa23):

ok I see. when you put it into the double angle you use the value from the triangle which we found using \[\tan ^{-1}\] correct ?

OpenStudy (phi):

yes. You found those correctly.

OpenStudy (saraa23):

Ok now that I have those values. how could I turn all my equation into an algebraically expression?

OpenStudy (phi):

\[ \cos 2( \tan^{-1} x) = \cos(2A) = \left(\cos A\right)^2 - \left(\sin A\right)^2 \]

OpenStudy (phi):

*\[ \cos(2 \tan^{-1}x) \]

OpenStudy (phi):

we know \[ \cos A= \frac{1}{\sqrt{x^2+1}}\] so plot that into the formula ditto for sin A

OpenStudy (phi):

you get \[ \frac{1}{x^2+1}- \frac{x^2}{x^2+1}\]

OpenStudy (saraa23):

so technically you just drop the cos and sin and keep the values we found using our triangle ?

OpenStudy (phi):

I'm not sure what you mean by drop the cosine and sine we are using the info from the tan(A)= x/1 to find the cosA and the sinA for this problem, and using those "values" in the double angle formula

OpenStudy (phi):

we are doing the same thing as, for example, cos(60)= cos(2*30) = cos^2(30) - sin^2(30) and we use a calculator (or memorize) the numerical values for cos(30) and sin(30) In this problem, we don't have numbers, we have "expressions" but it's the same idea.

OpenStudy (saraa23):

Ahhh okay its making sense. However when the question asks for an algebraic expression it is telling us to give them an equation to find x without cosine and sine ? so how do we get rid of cosine and sine. to make it an algebraic expression. Maybe I'm over thinking it or approaching it the wrong way.

OpenStudy (phi):

As you know we can find cos A and sin A as fractions (see above) we use those values in the double angle formula, and we will get \[ \cos\left(2 \tan^{-1} x\right) = \frac{1-x^2}{1+x^2} \]

OpenStudy (phi):

we are not "finding x", we are finding the cos of the mess "in terms of x"

OpenStudy (phi):

we can test that formula. For example, when x=1 we expect to get 0 (the top 1-x^2 will be 0) now let's do it step by step: tan^-1 (1) = 45º 2*45 = 90 cos(90º) = 0 it works!

OpenStudy (saraa23):

Okay im still a little confused after we found our CosA and SinA how did we get rid of the sqrt's that came from the hypotenuse ?

OpenStudy (phi):

do you know the formula for cos(2A) ?

OpenStudy (saraa23):

yes its \[\cos ^{2}(A)-\sin ^{2}(A)\]

OpenStudy (phi):

and the first term is cosine squared so square the value (or expression if you like) you found for cos A what do you get ?

OpenStudy (phi):

or multiply \[ \frac{1}{\sqrt{1+x^2} }\cdot \frac{1}{\sqrt{1+x^2} }\] top times top and bottom times bottom

OpenStudy (saraa23):

Oh okay thats when the square roots get cancelled because squaring a sqrt cancel each other out !

OpenStudy (phi):

exactly

OpenStudy (saraa23):

Okay I see now ! Thank you so much !!!

OpenStudy (phi):

It is tricky, so you might have to go over the logic and try it on other problems.

OpenStudy (saraa23):

Yes it is very Tricky. Would it be the same process when we dont have a double angle inside ?

OpenStudy (phi):

if it were just cos( tan^-1 x) i.e. cos(A) it's the same idea but easier if it's cos(A/2) you can use the same idea: find cosA and sinA and use them in the half angle formula

OpenStudy (phi):

and I suppose you could do cos(3A) as long as you can rewrite that in terms of cosA and sin A and in all these cases I am assuming A = tan^-1 or sin^-1 or cos^-1 in each case, you can label a triangle with 2 sides and find the 3rd side, and then find the other trig values i.e. sin A cos A or tan A if you needed it.

OpenStudy (saraa23):

Okay. but in the situation where we would have for example Sin(Sin^1(x)) The answer would be x correct because it is the composite ?

OpenStudy (saraa23):

i meant \[Sin(Sin ^{-1}(x))\]

OpenStudy (phi):

yes, the answer is immediately x but we could do it by saying A= sin^-1 x sinA = x/1 and we want sin (sin^-1 x) which (replacing sin^-1 x with A ) is sin A and of course sin A = x/1 = x

OpenStudy (phi):

but if we wanted cos (sin^-1 x) we would have to do it as above: let A = sin^-1 x and so sin A = x/1 and find the other side : sqr( 1-x^2) and then we can find cos A

OpenStudy (saraa23):

I'm sorry for all the questions I'm just trying to fully understand since my exam is coming up. But in the example ummm... Lets say \[Sin(\cos ^{-1}(x))\] We would first find the values to our triangle correct ?

OpenStudy (saraa23):

which makes \[\cos ^{-1}(x)\] = \[\cos \theta =x \div1 \]

OpenStudy (phi):

I would write A= cos^-1 x then apply cos to both sides to get cos A = x write x as a fraction, so x/1 cos A= x/1 now label the sides of the triangle: adj is x , hyp 1 and we find the 3rd side

OpenStudy (saraa23):

if our hypotenuse is 1 and we have x as our adj. then our opp would as well be x ?

OpenStudy (phi):

no, you use pythagoras

OpenStudy (phi):

a^2 + b^2 = c^2 let b be x and c be 1 and solve for a

OpenStudy (saraa23):

which is \[a^2 + b^2 = c^2\] and since we have hypo. and adj it would be \[a^2 + x^2 = 1\]

OpenStudy (phi):

yes

OpenStudy (saraa23):

so after making our triangle we come down to our equation sin(x/1) correct ? since we replaced cos^-1

OpenStudy (phi):

no, you have a "bug" in your thinking to keep things clear, remember cos^-1 returns an angle. I always rename , for example, A= cos^-1 x and change sin(cos^-1 x) to sin(A) so I know I want the sin of A

OpenStudy (saraa23):

okay but at the end how could i rewrite it into an algebraic equation if i simply have Sin(A)

OpenStudy (phi):

meanwhile, I use A = cos^-1 x to write cosA = x/1 find the 3rd side of the triangle find sin A = sqr(1-x^2) and we are done.

OpenStudy (saraa23):

ohhhhh So you are just naming the side that we find with our triangle. since its not a double angle

OpenStudy (phi):

yes. if the problem was sin (2 cos^-1 x) i.e. sin(2A) we would have to find both sin A and cos A and use them in the double angle formula to find sin2A

OpenStudy (saraa23):

Okay. I see what i was doing ! With double angles we use the formulas and when its a simple angle we just find the measure of the triangle and find the corresponding

OpenStudy (phi):

yes.

OpenStudy (saraa23):

perfect thank you.

OpenStudy (phi):

yw